# Universal properties of the many-body Lanczos algorithm at finite size

### v1

@@ Acknowledgments
We provide here the technical details of the paragraph Analytical arguments. We first discuss the criterion for the convergence of the cumulative products of Eq. (9), and, specifically, we study
@@
$-\log\left(\prod^{n}_{n^{\prime}\text{ odd}}e^{-\Gamma_{n^{\prime}}}\right)=\sum^{n}_{n^{\prime}\text{ odd}}\Gamma_{n^{\prime}}\sim\frac{1}{2}\sum^{n}_{n^{\prime}=1}\frac{\alpha}{(n^{\prime})^{\beta}}.$ (24)
@@
Finally, for $10\leq\beta<$ 1 a crude estimation gives
[…]
@@
Supplemental Material to:
Universal properties of the many-body Lanczos algorithm at finite size
Luca Capizzi,1 Leonardo Mazza,1,2 and Sara Murciano1
1Université Paris-Saclay, CNRS, LPTMS, 91405, Orsay, France
2Institut Universitaire de France, 75005, Paris, France
@@
The focus of this work are the universal properties of the Lanczos coefficients $b_{n}$ at finite size $L$, which are produced as follows. The Lanczos algorithm (LA) starts with a seed operator $O_{0}=\mathcal{O}$, and then builds an orthonormal basis of operators $\{O_{n}\}$ by repeatedly acting with the Liouvillian $\mathcal{L}$ to $O_{0}$ while orthonormalizing against the one obtained at the step before. In formulas, this means that, at each iteration, we do
$\begin{split}O^{\prime}_{n}&=\mathcal{L}O_{n-1}-b_{n-1}O_{n-2},\\ O_{n}&=\frac{O^{\prime}_{n}}{b_{n}},\\ b_{n}&=||O^{\prime}_{n}||,\end{split}$ (S1)
@@
with $b_{0}=0$ and $||\mathcal{O}||=\sqrt{\langle\mathcal{O}^{\dagger}\mathcal{O}\rangle}$. This implies that at any iteration of the algorithm, only three elements of the basis $\{O_{n}\}$ must be stored in memory and permits the computation of several thousands of $b_{n}$ for chain lengths up to $L=14$. However, as we have emphasized in the main text, as we increase $n$, the LA may fail to produce perfectly orthogonal vectors (a discussion of this known fact can be found, for instance, in Ref. [S1]). This means that, unless we store more than 3 vectors, we cannot guarantee that our Krylov-basis $\{O_{n}\}$ is orthonormal. To correct for orthogonalization errors, we keep every computed state in memory and apply a Gram–Schmidt procedure to each newly generated vector. However, storing all $O_{n}$ requires a large amount of memory, and $L=13$ is the largest system size for which we can perform the extra Gram-Schmidt orthogonalization step.
@@
Figure S1: Time evolution of the autocorrelation $\langle\mathcal{O}(t)\mathcal{O}\rangle$ function for the Ising model in Eq. (3) of the main text. In the top panels, $\mathcal{O}=\sigma_{1}^{z}$, while in the bottom $\mathcal{O}=\sigma_{1}^{z}\sigma_{2}^{z}$. The different colors correspond to different methods: the blue symbols have been obtained using Exact Diagonalization (ED), the orange lines are based on the standard Lanczos algorithm (SA) described in Eqs. (S1), while the green symbols are obtained using the LA with the full Gram-Schmidt orthogonalization (FO). The LA has been truncated at $n=1000$ for all the autocorrelators at $L=10$, while for $L=14$ the truncation occurs at $n=2000$. The plots clearly show that both algorithms reproduce the same autocorrelations despite the numerical instabilities of the SA.
@@
Figure S2: Comparison between the cumulative products $F(n)$ obtained through the 3-step algorithm of Eq. (S1) (yellow) and with the full orthogonalization (green) for different system sizes $L=9,10,11,12$ (bottom to top). The dashed lines correspond to bi-exponential fits for the SA data (black) and the FO ones (red). Both panels show a good qualitative and quantitative agreement between the parameters extracted from the fit.
@@
Given these two methods, one can ask what is the impact of the numerical instabilities of the Lanczos coefficients in the quantities we are interested in here, namely the autocorrelation function in Eq. (1) of the main text, and the cumulative products in Eq. (15) of the main text. To address this question, we compare in Fig. S1 the behaviour of the autocorrelations of $\mathcal{O}=\sigma^{z}_{1}$ and $\sigma^{z}_{1}\sigma_{1}^{z}$ calculated by exact diagonalization (ED) with those computed using the LA both with (FO, full orthogonalization) and without (SA, standard algorithm) the full Gram-Schmidt orthogonalization. With FO, the maximum value of $n$ we can reach for $L=10$ is $n=1000$. For this reason, we restrict to the same window for the three-step standard algorithm of Eq. (S1) for $L=10$, while we reach $n=2000$ for $L=14$. In the three plots of Fig. S1, we clearly observe that the discrepancies arising from the imperfect orthogonalization procedure does not have any strong impact on the time-evolution of the autocorrelations, which is the main focus of this work. The deviations from the exact results are $O(10^{-4})$ with both approaches, also at larger system sizes.
@@
Figure S3: Behavior of $-\log F(n)\cdot L^{m+1}/n$ for the operators $\mathcal{O}=\sigma_{1}^{z}\sigma_{2}^{z}$ ($m=2$, left panel) and $\mathcal{O}=\sigma_{1}^{z}\sigma_{2}^{z}\sigma_{3}^{z}$ ($m=3$, right panel). Using $6000$ Lanczos coefficients $b_{n}$, the different curves collapse at large values of $n$, confirming the validity of our first conjecture, summarized by Eq. (6) in the main text.
In Fig. S2, we also apply the bi-exponential fitting of $F(n)\sim a_{1}e^{-c_{1}n}+a_{2}e^{-c_{2}n}$ for the data with and without full orthogonalization. Remarkably, the fit function works very well in both cases and the fit parameters extracted from the Lanczos coefficients computed with full Gram–Schmidt orthogonalization are close to those obtained using the three-step algorithm in Eq. (S1), yielding a similar average rate $\bar{\Gamma}$. Not only, thus, the data affected by the numerical error collapse according to the same qualitative function, they also display almost identical quantitative features.
Given the similarities of the data produced by the two algorithms, with and without full orthogonalization, we check the scaling rate of Eq. (6) in the main text for $m=2,3$ with the operators $\mathcal{O}=\sigma_{1}^{z}\sigma_{2}^{z}$ and $\mathcal{O}=\sigma_{1}^{z}\sigma_{2}^{z}\sigma_{3}^{z}$, respectively, using the SA in Eq. (S1) in Fig. S3. This allows us to reach larger values of $n\sim 6000$. These data are out of reach if one considers the LA with full orthogonalization. We notice that a slow collapse of the data is visible only at values of $n$ which are larger than those considered in Fig. 3 of the main text for $\mathcal{O}=\sigma_{1}^{z}$. This supports our conjecture; the reason why the collapse was not visible for the operators $\sigma^{z}_{1}\sigma^{z}_{2}$ and $\sigma^{z}_{1}\sigma^{z}_{2}\sigma^{z}_{3}$ using the LA with full orthogonalization (and was not presented in the main text) is likely that we could not compute the $b_{n}$ for large-enough $n$. In summary, the results presented in this Supplemental Material suggest that the standard Lanczos algorithm gives access to the universal features of the finite-size Lanczos coefficients $b_{n}$ no matter its numerical instability.
We conclude this Supplementary Material by examining the origin of the imperfect data collapse in our plots of
$-\log F(n)\cdot L^{m+1}/n\text{ for }m=1,2$ (S2)
shown in Fig. 3 of the main text and Fig. S3, respectively. Our results rely on the assumption that the plateaus $C_{L}(\infty)$ scale as $L^{-(m+1)}$. For the system sizes we have access to using exact diagonalization, we show in Fig. S4 that the plateau of the autocorrelations of $\sigma^{z}_{1}$ (left panel) and $\sigma^{z}_{1}\sigma^{z}_{2}$ (right panel) decays faster than $L^{-1}$ or $L^{-2}$, respectively (especially in the latter case), due to finite size effects. This explains why the collapse of the data of the cumulative products is not perfect. We are convinced that if we could examine larger systems, we could observe the scalings predicted in the main text.
@@
Figure S4: Autocorrelation function of $\mathcal{O}=\sigma_{1}^{z}$ (left panel) and $\mathcal{O}=\sigma_{1}^{z}\sigma_{2}^{z}$ (right panel) using exact diagonalization. The insets are a fit of the plateau for different system sizes. While in the left inset the data for larger systems approach the scaling $L^{-1}$ of the orange dashed line, in the right inset the deviations from the reference line $L^{-2}$ are more evident.
Bibliography of the Supplemental Material
  1. [S1]
    E. Rabinovici, A. Sánchez-Garrido, R. Shir, J. Sonner, Operator complexity: a journey to the edge of Krylov space, J. High Energ. Phys. 2021, 62 (2021). https://doi.org/10.1007/JHEP06(2021)062

### v2

@@ IV.1 Technical details
We provide here the technical details about the cumulative products in Eq. (17). We first discuss the criterion for the convergence of the cumulative products of Eq. (16), and, specifically, we study
@@
$\log\left(\prod^{n}_{n^{\prime}\text{ odd}}e^{-\Gamma_{n^{\prime}}}\right)=\sum^{n}_{n^{\prime}\text{ odd}}\Gamma_{n^{\prime}}\sim\frac{1}{2}\sum^{n}_{n^{\prime}=1}\frac{\alpha}{(n^{\prime})^{\beta}}.$ (22)
@@
Finally, for $0\leq\beta<1$ a crude estimation gives
$\frac{1}{2}\sum^{n}_{n^{\prime}=1}\frac{\alpha}{(n^{\prime})^{\beta}}\sim\frac{\alpha}{2}\int^{n}_{1}dx\frac{1}{x^{\beta}}\propto\alpha n^{1-\beta},$ (24)
@@
V Numerics
We now present supporting numerical results, obtained with an exact-diagonalization algorithm using up to 600 GB of RAM; since the computation of the $b_{n}$ coefficients with the LA is numerically unstable, we resort to the full Gram-Schmidt orthogonalization scheme, which is computationally more expensive. It is a remarkable fact discussed in Appendix A that the coefficients computed with the standard LA algorithm describe the same autocorrelation function and satisfy the same universal scaling laws that we present here below.
V.1 Simple checks: short-range one-dimensional models
In this section, we focus on the model in Eq. (3); additional numerical studies addressing long-range systems and higher-dimensional lattices are presented in Sec. V.2, supporting our conjecture and its claimed universality. We consider the three observables $\mathcal{O}=\sigma^{z}_{1}$, $\sigma^{z}_{1}\sigma^{z}_{2}$ and $\sigma^{y}_{1}$. According to the analysis presented in Ref. [48], $\sigma^{z}_{1}$ overlaps (in the infinite temperature state) with $H$ and $\sigma^{z}_{1}\sigma^{z}_{2}$ with $H^{2}$: as a consequence, the plateaus $C_{L}(\infty)$ scale respectively as $L^{-1}$ and $L^{-2}$. Hence, according to the Conjecture 1, $\Gamma_{n}(L)$ is expected to approach a positive constant $\sim L^{-(m+1)}$, with $m=1,2$ respectively.
Figure 2: The cumulative product $F(n)$ for the operators $\sigma^{z}_{1}$ (left panel) and $\sigma^{z}_{1}\sigma^{z}_{2}$ (right panel). The data refer to the Ising Hamiltonian (3) with parameters $J,h_{x},h_{z}]=[1,1,1.5]$, and the dashed black lines show the fit with bi-exponential curves.
Figure 3: The cumulative products in Fig. 3 for the operator $\sigma^{z}_{1}$ ($m=1$) rescaled as $\log F(n)\cdot L^{m+1}/n$. Assuming conjecture (6), the curves are expected to collapse for large $L$.
$F(n):=\prod^{n}_{n^{\prime}=0}e^{-\Gamma_{n^{*}+2n^{\prime}}(L)}.$ (27)
For both operators, the data are consistent with the presence of a decay rate that decreases with the system size. In order to check whether the scaling of the rate with $L$ is compatible with Eq. (6), in Fig. 3 we plot the quantity $\log F(n)\cdot L^{m+1}/n$ as a function of $n$ for different sizes: these curves are expected to collapse for large $L$ toward a constant value. The results for $\sigma^{z}_{1}$ in Fig. 3 display a good agreement with the conjecture; the data for $\sigma^{z}_{1}\sigma^{z}_{2}$ are less convincing, and a better agreement would be found if larger values of $n$ were accessible: we comment on the issue in the Appendix A.
We now move to the study of the operator $\sigma^{y}_{1}$, which does not overlap with any power of $H$ (being odd under time reversal) [48]; the infinite-time plateau value of $C_{L}(\infty)$ is expected to vanish and thus it allows to test Conjecture 2. Our numerics is presented in Fig. 4 (left panel), where we show that in this case $F(n)$ is an increasing function of $n$. This means that $\Gamma_{n}(L)$ fluctuates around a negative value, and thus is compatible with the inequality (7), which summarizes Conjecture 2.
Finally, we consider the Hamiltonian
with open boundary conditions and parameters $\mu=0.2$, $U=0.5$; the model is known to host a boundary zero mode at zero temperature that becomes approximate when the temperature is increased [16]. We probe the operator $\mathcal{O}=\sigma^{x}_{1}$ at the left boundary of the chain in the infinite temperature state. We show the results in Fig. 4 (right panel): the data for $F(n)$ are compatible with a saturation at large $n$, consistent with $\Gamma_{n}\simeq 0$, implying $C_{L}(\infty)=0$ according to Eq. (18). If the zero mode were exact, we would have observed an exponential decay of $F(n)$ with a size-independent decay rate $\Gamma$. Our data are thus compatible with Conjecture 2, although from these numerical data it is difficult to extract reliable quantitative values for the large $n$ scalings of $\Gamma_{n}(L)$. By comparing the two panels of Fig. 4, we observe a marked qualitative difference between the two cases, which share the common property $C_{L}(\infty)=0$, but are associated to two qualitatively distinct physical situations (absence of hydrodynamic projection versus approximate zero mode).
V.2 Further numerical benchmarks: long-range and higher dimensional systems
We consider a one-dimensional Ising model with long-range interactions:
$H=\sum_{i=1}^{L}\sum_{j>i}\frac{1}{N_{\alpha}r_{ij}^{\alpha}}\sigma^{x}_{i}\sigma^{x}_{j}+\sum_{i=1}^{L}h_{z}\sigma^{z}_{i}$ (29)
where
$r_{ij}=\mathrm{min}(|j-i|,L-|j-i|),\quad N_{\alpha}=\left(\sum_{i=2}^{L}\frac{1}{r_{i1}^{2\alpha}}\right)^{1/2},$ (30)
and we focus on the operator $\mathcal{O}=\sigma^{z}_{1}$. The computed values of $b_{n}$ are shown in Fig. 5 (left). Here, since this model is long-ranged, the values of $b_{n}$ do not coincide exactly across different sizes for sufficiently small $n$, unlike in short-range systems (see, for instance, Fig. 1). Nonetheless, a qualitative agreement is observed, and for $n>n^{*}$ the Lanczos coefficients saturate to a plateau. Qualitatively, thus, we have the same behavior displayed in Fig. 1.
We then analyse the $b_{n}$ coefficients and in Fig. 5 (right) we plot $F(n)$ in Eq. (27), choosing $n^{*}$ with the same convention as in the short-range case, and fit the data with a bi-exponential, obtaining an excellent agreement. As in the short-range model, the long-range system shows an asymptotic exponential decay of $F(n)$, with a decay rate that diminishes as $L$ increases.
However, in the presence of long-range interactions, we have no reason to expect $n^{*}$ to scale as $\sim L$, and therefore we cannot directly employ the prediction of Eq. (6). To the best of our knowledge, this problem has never been studied theoretically in the literature, and the accessible sizes are not sufficiently large to propose a convincing ansatz for the behavior of $n^{*}$ in long-range models based on numerical data. Nonetheless, following the steps leading to Eq. (21) and assuming $b_{n}$ is linear in the initial regime, we deduce
$C_{L}(\infty)\sim n^{*}\Gamma(L).$ (31)
Thus, if $C_{L}(\infty)\sim 1/L^{m}$ (here we expect $m=1$, even if the model is long-range; see Ref. [48] for details) and $n^{*}\sim L^{a}$ for some positive value of $a$, we can still infer an algebraic decay of $\Gamma(L)$ as
$\Gamma(L)\sim L^{-m-a}.$ (32)
In Fig. 6, we test the compatibility of our data with an algebraic decay $\Gamma(L)$ by plotting the function $\log F(n)L^{1+a}/n$. In the left panel, we consider $a=1$, which corresponds to the scaling expected from Eq. (6), and namely with Conjecture 1. In the right panel, we choose $a=0.5$, a value that exhibits better compatibility with a collapse. Although we cannot make strong claims about the precise value of $a$, it is reasonable to expect $a<1$, since long-range interactions tend to accelerate the spreading of operators, thereby decreasing $n^{*}$ compared to their short-range counterparts.
We also consider a two-dimensional short-range model
$H=\sum_{\langle\mathbf{i},\mathbf{j}\rangle}J\sigma_{\mathbf{i}}^{x}\sigma_{\mathbf{j}}^{x}+\sum_{\mathbf{j}}(h_{x}\sigma_{\mathbf{j}}^{x}+h_{z}\sigma_{\mathbf{j}}^{z}),$ (33)
where $\langle\mathbf{i},\mathbf{j}\rangle$ denotes the nearest neighbours on a $L_{x}\times L_{y}$ lattice with open boundary conditions. We consider $(L_{x},L_{y})=(3,3),(3,4)$ and we focus on the operator $\sigma^{z}_{(2,2)}$ (inserted at the center of the $3\times 3$ lattice). In Fig. 7 (left), we show the Lanczos coefficients $b_{n}$: they display an initial linear growth, $b_{n}\sim n$, and a plateau for $n>n^{*}$, resembling the behaviour observed for one-dimensional systems. Similarly, we show the decay of the function $F(n)$ in Fig. 7 (right), and the numerical data are compatible with a bi-exponential decay.
In this case, the computational cost limits access to larger system sizes with standard numerics, making the finite-size scaling analysis difficult. Nonetheless, since the value of the plateau is expected to scale as $1/L^{d}$ ($d=2$, and $L$ the typical size of the lattice) [48], and the expected value of $n^{*}$ is $n^{*}\sim L$, as a consequence of locality of the interactions, we predict $\Gamma(L)\sim L^{-3}$ compatible with the Conjecture 1 and specifically with Eq. (6). This is consistent with numerical data, since we observe that the rate $\Gamma_{n}(L)$ is smaller when the lattice is bigger.
Figure 5: (Left) Lanczos coefficients $b_{n}$ associated with the operator $\sigma^{z}_{1}$ for the long-range Ising Hamiltonian in Eq. (29) with $\alpha=1.5,h_{z}=1.5$. (Right) The function $F(n)$ for different system sizes; the colored dots are numerical data, while the dashed black lines are biexponential fits.
Figure 6: We plot the quantity $L^{1+a}\log F(n)/n$ for the operator $\sigma^{z}_{1}$: we consider $a=1$ (left) and $a=0.5$ (right). The data refer to the long-range Ising Hamiltonian in Eq. (29) with $\alpha=1.5,h_{z}=1.5$.
Figure 7: We plot the Lanczos coefficients $b_{n}$ (left panel) and $F(n)$ (right panel) for the operator $\mathcal{O}=\sigma^{z}_{(2,2)}$. The data refer to 2D Ising Hamiltonian in Eq. (33) with parameters $J,h_{x},h_{z}]=[1,1,1.5]$ and lattice sizes $(L_{x},L_{y})=(3,3),(3,4)$.
VI Conclusions and outlook
We considered the LA for quantum many-body systems, and we focused on the Lanczos coefficients $b_{n}$, computing several hundred of them for finite-size systems. The universal properties of autocorrelation functions in generic models allowed us to present a threefold conjecture on the scaling of the ratios $(b_{n}/b_{n+1})^{2}$, with $n$ odd, in the large $n$ limit. Our main goal was to show that the finite-size behavior of the Lanczos coefficients has universal properties that can be studied with state-of-the-art numerics.
This work opens intriguing research directions. A natural question is to understand the origin of the bi-exponential behavior found in Fig. 3. Moreover, after having studied the late-time plateau of $C_{L}(t)$, it is natural to focus on the algebraic decay $t^{-\nu}$ (partially addressed in Ref. [47]), and to investigate how the finite-size Lanczos coefficients give rise to such a behavior. Also, the precise mechanism by which the finite-size crossover of $C_{L}(t)$, from decay to plateau, occurs at the level of the LA is currently unclear. Finally, our study of the Lanczos coefficients has highlighted an erratic behavior in the region $n>n^{*}$; whether the latter has some universal properties, and it can be studied as a stochastic process, is a question that deserves to be addressed.
Acknowledgements.
@@
The focus of this work is the universal properties of the Lanczos coefficients $b_{n}$ at finite size $L$, which are produced as follows. The Lanczos algorithm (LA) starts with a seed operator $O_{0}=\mathcal{O}$, and then builds an orthonormal basis of operators $\{O_{n}\}$ by repeatedly acting with the Liouvillian $\mathcal{L}$ to $O_{0}$ while orthonormalizing against the one obtained at the step before. In formulas, this means that, at each iteration, we compute
[…]
@@
with $b_{0}=0$ and $\mathcal{O}||=\sqrt{\langle\mathcal{O}^{\dagger}\mathcal{O}\rangle}$. This implies that at any iteration of the algorithm, only three elements of the basis $\{O_{n}\}$ must be stored in memory and permits the computation of several thousands of $b_{n}$ for chain lengths up to $L=14$. However, as we have emphasized in the main text, as we increase $n$, the LA may fail to produce perfectly orthogonal vectors (a discussion of this known fact can be found, for instance, in Ref. [22]). This means that, unless we store more than 3 vectors, we cannot guarantee that our Krylov-basis $\{O_{n}\}$ is orthonormal. To correct for orthogonalization errors, we keep every computed state in memory and apply a Gram–Schmidt procedure to each newly generated vector. However, storing all $O_{n}$ requires a large amount of memory, and $L=13$ is the largest system size for which we can perform the extra Gram-Schmidt orthogonalization step.
@@
Figure 8: Time evolution of the autocorrelation $\langle\mathcal{O}(t)\mathcal{O}\rangle$ function for the Ising model in Eq. (3) of the main text. In the top and central panels, $\mathcal{O}=\sigma_{1}^{z}$, while in the bottom $\mathcal{O}=\sigma_{1}^{z}\sigma_{2}^{z}$. The different colors correspond to different methods: the blue symbols have been obtained using Exact Diagonalization (ED), the orange lines are based on the standard Lanczos algorithm (SA) described in Eqs. (34), while the green symbols are obtained using the LA with the full Gram-Schmidt orthogonalization (FO). The LA has been truncated at $n=1000$ for all the autocorrelators at $L=10$, while for $L=14$ the truncation occurs at $n=2000$. The plots clearly show that both algorithms reproduce the same autocorrelations despite the numerical instabilities of the SA.
@@
Figure 9: Comparison between the cumulative products $F(n)$ obtained through the 3-step algorithm of Eq. (34) (yellow) and with the full orthogonalization (green) for different system sizes $L=9,10,11,12$ (bottom to top). The dashed lines correspond to bi-exponential fits for the SA data (black) and the FO ones (red). Both panels show a good qualitative and quantitative agreement between the parameters extracted from the fit.
@@
Given these two methods, one can ask what is the impact of the numerical instabilities of the Lanczos coefficients in the quantities we are interested in here, namely the autocorrelation function in Eq. (1) of the main text, and the cumulative products in Eq. (15) of the main text. To address this question, we compare in Fig. 8 the behaviour of the autocorrelations of $\mathcal{O}=\sigma^{z}_{1}$ and $\sigma^{z}_{1}\sigma_{1}^{z}$ calculated by exact diagonalization (ED) with those computed using the LA both with (FO, full orthogonalization) and without (SA, standard algorithm) the full Gram-Schmidt orthogonalization. With FO, the maximum value of $n$ we can reach for $L=10$ is $n=1000$. For this reason, we restrict to the same window for the three-step standard algorithm of Eq. (34) for $L=10$, while we reach $n=2000$ for $L=14$. In the three plots of Fig. 8, we clearly observe that the discrepancies arising from the imperfect orthogonalization procedure does not have any strong impact on the time-evolution of the autocorrelations, which is the main focus of this work. The deviations from the exact results are $O(10^{-4})$ with both approaches, also at larger system sizes.
@@
Figure 10: Behavior of $\log F(n)\cdot L^{m+1}/n$ for the operators $\mathcal{O}=\sigma_{1}^{z}\sigma_{2}^{z}$ ($m=2$, top panel) and $\mathcal{O}=\sigma_{1}^{z}\sigma_{2}^{z}\sigma_{3}^{z}$ ($m=3$, bottom panel). Using 60006000 Lanczos coefficients $b_{n}$, the different curves collapse at large values of $n$, confirming the validity of our first conjecture, summarized by Eq. (6) in the main text.
@@
Figure 11: Autocorrelation function of $\mathcal{O}=\sigma_{1}^{z}$ (top panel) and $\mathcal{O}=\sigma_{1}^{z}\sigma_{2}^{z}$ (bottom panel) using exact diagonalization. The insets are a fit of the plateau for different system sizes. While in the left inset the data for larger systems approach the scaling $L^{-1}$ of the orange dashed line, in the right inset the deviations from the reference line $L^{-2}$ are more evident.
In Fig. 9, we also apply the bi-exponential fitting of $F(n)\sim a_{1}e^{-c_{1}n}+a_{2}e^{-c_{2}n}$ for the data with and without full orthogonalization. Remarkably, the fit function works very well in both cases and the fit parameters extracted from the Lanczos coefficients computed with full Gram–Schmidt orthogonalization are close to those obtained using the three-step algorithm in Eq. (34), yielding a similar average rate $\bar{\Gamma}$. Not only, thus, the data affected by the numerical error collapse according to the same qualitative function, they also display almost identical quantitative features.
Given the similarities of the data produced by the two algorithms, with and without full orthogonalization, we check the scaling rate of Eq. (6) in the main text for $m=2,3$ with the operators $\mathcal{O}=\sigma_{1}^{z}\sigma_{2}^{z}$ and $\mathcal{O}=\sigma_{1}^{z}\sigma_{2}^{z}\sigma_{3}^{z}$, respectively, using the SA in Eq. (34) in Fig. 10. This allows us to reach larger values of $n\sim 6000$. These data are out of reach if one considers the LA with full orthogonalization. We notice that a slow collapse of the data is visible only at values of $n$ which are larger than those considered in Fig. 3 of the main text for $\mathcal{O}=\sigma_{1}^{z}$. This supports our conjecture; the reason why the collapse was not visible for the operators $\sigma^{z}_{1}\sigma^{z}_{2}$ and $\sigma^{z}_{1}\sigma^{z}_{2}\sigma^{z}_{3}$ using the LA with full orthogonalization (and was not presented in the main text) is likely that we could not compute the $b_{n}$ for large-enough $n$. In summary, the results presented in this Appendix suggest that the standard Lanczos algorithm gives access to the universal features of the finite-size Lanczos coefficients $b_{n}$ no matter its numerical instability.
We conclude this Appendix by examining the origin of the imperfect data collapse in our plots of $\log F(n)\cdot L^{m+1}/n(m=1,2)$ shown in Fig. 3 of the main text and Fig. 10, respectively. Our results rely on the assumption that the plateaus $C_{L}(\infty)$ scale as $L^{-(m+1)}$. For the system sizes we have access to using exact diagonalization, we show in Fig. 11 that the plateau of the autocorrelations of $\sigma^{z}_{1}$ (top panel) and $\sigma^{z}_{1}\sigma^{z}_{2}$ (bottom panel) decays faster than $L^{-1}$ or $L^{-2}$, respectively (especially in the latter case), due to finite size effects. This explains why the collapse of the data of the cumulative products is not perfect. We are convinced that if we could examine larger systems, we could observe the scalings predicted in the main text.